# ControlFace: Harnessing Facial Parametric Control for Face Rigging

### v1

@@ 4.3 Injecting Appearance of Reference Image
To acquire strong image representations that preserve identity and fine details during generation, the reference image is encoded using both the CLIP [32] image encoder and the proposed FaceNet. The CLIP image embeddings, which capture high-level semantic information, are integrated into both denoising U-Net and FaceNet through cross-attention layers. Simultaneously, FaceNet encodes the reference facial image to extract detailed features which are then incorporated into the denoising U-Net through augmented self-attention mechanism [19, 48, 38]. Since FaceNet and the denoising U-Net share the same architecture, they integrate seamlessly in self-attention by concatenating keys and values from both U-Nets and applying attention operation with queries from the denoising U-Net. Let $Q$, $K$, $V\in\mathbb{R}^{l\text{x}d}$ denote query, key, value from the denoising U-Net where $l$ and $d$ denote length and dimensionality. Also, let $K^{\mathrm{face}}$, $V^{\mathrm{face}}\in\mathbb{R}^{l\text{x}d}$ represent key, and value from the FaceNet, respectively. Then the augmented self-attention $\mathrm{Aug\text{-}Attn}$ is formulated as follows:
@@ 4.5 Training Objective
Since we perform the generation process in the compact latent space of a VAE [22], we compress $X_{R}$ via the VAE encoder and acquire reference latent $z_{R}$. Our model $\epsilon_{\theta}(\cdot)$ takes $z_{R}$, $D_{T}$, and $D_{R}$ as input and learns to predict random Gaussian Noise $\epsilon$ which is used to perturb $z_{R}$. Similar to Eq. 1, our training objective is as follows:
$\mathcal{L}=\mathbb{E}_{z,t,\epsilon\sim\mathcal{N}(0,1)}[\|\epsilon_{\theta}(z_{R,t};t,z_{R},D_{T},D_{R})-\epsilon\|^{2}_{2}],$ (5)
@@ 5.3 Result on Out-of-Domain images
Figure 8: Ablation on Guidance Scale. We plot the DECA [10] re-inference error with varying guidance scale $w$ from 1.0 to 4.0. As the scale increases, the DECA re-inference error improves for RCF and $\text{CFG}^{*}$, while worsens for DiffusionRig [8].
@@ 5.4 Ablation Studies
We examine how varying guidance scales $w$ affects control adherence. From here, $\text{CFG}^{*}$ denotes CFG applied to the face controller input whereas CFG means applying null input to the CLIP encoder similar to most text-to-image models [35, 9, 34]. Since RCG requires no null condition training, we also apply it to DiffusionRig, evaluating on 128 randomly selected FFHQ reference images. Fig. 4 shows that increasing the scale improves DECA re-inference error for RCG and $\text{CFG}^{*}$ but reduces performance for DiffusionRig. We suspect this is due to DiffusionRig’s struggle to maintain identity consistency under different controls, causing conflicting identities and poor alignment when using both target and reference controls.
@@
Starting with FaceNet and Face Controller, we sequentially add CMM, CFG [17], $\text{CFG}^{*}$, and our proposed RCG, with results in Tab. 5. Incorporating CMM improves alignment and image quality but slightly reduces identity similarity compared to FaceNet. This occurs because enhanced control adherence requires adjustments to attributes like expression, pose, or lighting, which may inadvertently alter identity-related features. Consequently, FaceNet, with poorer alignment, achieves the highest identity preservation. Our proposed RCG method compensates for the marginal decrease in ID score by achieving better control adherence and image quality than FaceNet. Also, RCG exhibits better control alignment among other guidance methods without sacrificing the identity. We additionally provide ablation study on different inputs for CCM in the Appendix C.3.

### v4

@@ 4.3 Injecting Appearance of Reference Image
To acquire strong image representations that preserve identity and fine details during generation, the reference image is encoded using both the CLIP [32] image encoder and the proposed FaceNet. The CLIP image embeddings, which capture high-level semantic information, are integrated into both denoising U-Net and FaceNet through cross-attention layers. Simultaneously, FaceNet encodes the reference facial image to extract detailed features which are then incorporated into the denoising U-Net through augmented self-attention mechanism [19, 49, 38, 41]. Since FaceNet and the denoising U-Net share the same architecture, they integrate seamlessly in self-attention by concatenating keys and values from both U-Nets and applying attention operation with queries from the denoising U-Net. Let $Q$, $K$, $V\in\mathbb{R}^{l\text{x}d}$ denote query, key, value from the denoising U-Net where $l$ and $d$ denote length and dimensionality. Also, let $K^{\mathrm{face}}$, $V^{\mathrm{face}}\in\mathbb{R}^{l\text{x}d}$ represent key, and value from the FaceNet, respectively. Then the augmented self-attention $\mathrm{Aug\text{-}Attn}$ is formulated as follows:
@@ 4.5 Training Objective
Since we perform the generation process in the latent space of a VAE [22], we encode $X_{R}$ and $X_{T}$ via the VAE encoder and acquire latents $z_{R}$ and $z_{T}$, respectively. Our model $\epsilon_{\theta}(\cdot)$ takes $z_{T}$,$z_{R}$, $D_{T}$, and $D_{R}$ as input and learns to predict random Gaussian Noise $\epsilon$ which is used to perturb $z_{R}$. Similar to Eq. 1, our training objective is as follows:
$\mathcal{L}=\mathbb{E}_{z,t,\epsilon\sim\mathcal{N}(0,1)}[\|\epsilon_{\theta}(z_{T,t};t,z_{R},D_{T},D_{R})-\epsilon\|^{2}_{2}],$ (5)
@@ 5.3 Result on Out-of-Domain images
Figure 8: Ablation on Guidance Scale. We plot the DECA [10] re-inference error for guidance scales $w$ from 1.0 to 4.0. Increasing $w$ improves errors for RCG, $\text{CFG}^{*}$, and $\text{CFG}^{\dagger}$ to some extents, but greatly worsens for DiffusionRig [8].
@@ 5.4 Ablation Studies
We examine the effect of guidance scales $w$ on control adherence. From here, $\text{CFG}^{*}$ and $\text{CFG}^{\dagger}$ refer to CFG applied to the face controller and CMM, respectively, both incorporating null inputs. On the other hand, CFG represents applying a null input to the CLIP encoder, following the approach used in most text-to-image models [35, 9, 34]. Since RCG requires no null condition training, we also apply it to DiffusionRig, evaluating on 128 randomly selected FFHQ reference images. Fig. 8 shows that increasing the scale improves DECA re-inference error for RCG, $\text{CFG}^{*}$, and $\text{CFG}^{\dagger}$ to varying extents, but significantly reduces performance for DiffusionRig. We suspect this is due to DiffusionRig’s struggle to maintain identity consistency under different controls, causing conflicting identities and poor alignment when using both target and reference controls.
@@
Starting with FaceNet and Face Controller, we sequentially add CMM, CFG [17], $\text{CFG}^{*}$, and our proposed RCG, with results in Tab. 5. Incorporating CMM improves alignment and image quality but slightly reduces identity similarity compared to FaceNet. This occurs because enhanced control adherence requires adjustments to attributes like expression, pose, or lighting, which may inadvertently alter identity-related features. Consequently, FaceNet, with poorer alignment, achieves the highest identity preservation. Our proposed RCG method compensates for the marginal decrease in ID score by achieving better control adherence and image quality than FaceNet. Also, RCG exhibits better control alignment among other guidance methods without sacrificing the identity. We additionally provide ablation study on different inputs and architecture for CCM in the Appendix C.3.
@@ C.3 More ablation on CMM
We also trained three models with different architectures on CMM. Fig. 10 shows that our design of CMM exhibits the best results.
@@
Table 10: Different architectures for CMM. We train the model with different architectures for the CMM and measure the DECA [10] re-inference error and identity similarity (ID) [5].
Variants Re-Infer ($\downarrow$) ID ($\uparrow$)
CMM with ResNet architecture 5.27 0.7199
No CMM, Face Controller on reference 5.40 0.7184
CMM (Reference only) 5.03 0.7290
CMM (Ours) 4.85 0.7586